If cos(A-B)=1/2 and sin(A+B)=1/2, find the smallest values of A and B and also their general values?

Grade:12

1 Answers

Likith Narayan ss
12 years ago

cos(A-B)=1/2=>cos(A-B)=cos60    

                      =>A-B=60            --->1     

sin(A+B)=1/2=>sin(A+B)=sin30                

                      =>A+B=30             --->2                

comparing them we get A=45 B=15(in clockwise direction)